please help with this problem:
Bob fenced in a rectangular garden in his yard. The length of the rectangle is 5 feet longer than the width and the perimeter is 70 feet. What is the width of the rectangle?
thanks

Answer by mathwhiz(13)

70 / 2 = 35
5 / 2 = 2.5
35 - 2.5 = 32.5(width)
35 + 2.5 = 37.5(length)
37.5 - 32.5 = 5
The width of the rectangle is 32.5 feet.